State the pressure law and describe an experiment to verify it. Deduce the formula for the pressure coefficient of gas. What is this value?
1
Expert's answer
2020-06-01T14:28:16-0400

The pressure law: At a constant volume, if the mass of the gas does not change, the pressure over temperature ratio is a constant value:


"\frac{P_1}{T_1}=\frac{P_2}{T_2}=\text{const}."

Deduce the formula for the pressure coefficient of gas. Start with the ideal gas law:


"PV=nRT,\\
P=\frac{nR}{V}\cdot T=kT,\\
k=\frac{nR}{V}."

The value depends on the volume of gas and the amount of the substance.